OpenStudy (babynini):

Find the horizontal and vertical components of the vector with the given length and direction, and write the vector in terms of the vectors i and j. |v| = 3, θ = 225°

OpenStudy (babynini):

I got that this equals v<[(-3sq2)/2]i,[(-3sq2)/2]j>

OpenStudy (loser66):

yup, but i, j form should be ....i +.....j However, depends on your prof's note. You are good.

OpenStudy (babynini):

it says it's wrong o.0

OpenStudy (babynini):

instead of a comma it should be a + ?

OpenStudy (loser66):

ok, try \(\vec v= \dfrac{-3\sqrt2}{2}i -\dfrac{3\sqrt 2}{2}j\)

OpenStudy (babynini):

nopee doesn't want it.

OpenStudy (loser66):

:) good luck.

OpenStudy (babynini):

oh gosh ahha thanks.

OpenStudy (babynini):

on the left is how i entered my answer, on the right is how it says to enter the answer ( the example problem)

OpenStudy (babynini):

@Loser66

OpenStudy (babynini):

I'll just ask the prof later -.- thanks for your help!